1)twice the sum of a number and eleven is twenty-two less than three times the nmber. find the number.
2)the sum of three consecutive even integers is 66 find the integers.
3) the perimeter of a triangle is 96 in. The lengths of the sides are represented by consecutive integars. find the measures of the sides.
4)the sum of three consecutive odd integars is 89 more than twice the largest integar. find the integars.
thank you for your help

Answer by Alan3354(69443)

1)twice the sum of a number and eleven is twenty-two less than three times the nmber. find the number.
2n + 11 = 3n - 22
n = 33
-------
2)the sum of three consecutive even integers is 66 find the integers.
Do these the easy way.
66/3 = 22, that's the middle number.
--> 20, 22, 24
------------
3) the perimeter of a triangle is 96 in. The lengths of the sides are represented by consecutive integars. find the measures of the sides.
Same thing as above.
96/3 = 32
Sides = 31, 32, 33
------------------
4)the sum of three consecutive odd integars is 89 more than twice the largest integar. find the integars.
n = middle integer
n-4 + n-2 + n = 2n + 89
3n-6 = 2n+89
n = 95
--> 91, 93, 95
-----
PS, it's integer, no a.